# TS-Haystack: A Multi-Scale Retrieval Benchmark for Time Series Language Models

### v1

@@ 2 Background
Approaches such as Softprompting or Time-LLM (Jin et al., 2024), map the sequence of $N$ patch embeddings to the language space. However, this fidelity comes at a prohibitive computational cost ($\mathcal{O}(N^{2})$) either through the LLM’s extended context window or within the time series encoder, which must perform self-attention over the full patch sequence before projection (Wang et al., 2025). Consequently, these methods struggle to scale to larger context lengths due to computational cost and large memory requirements.
@@ B.7 Task Details
Construction: (1) Sample background window; (2) Select activity regime (sedentary/active); (3) Insert $N$ needles from regime activities not in background; (4) For positive samples, ask about an inserted activity; for negative, ask about a non-inserted activity from the same regime.
@@
Construction: (1) Sample two distinct activities $A$ and $B$; (2) Sample background excluding both; (3) Randomly assign temporal order (50/50 balance); (4) Insert both needles sequentially.
@@
Construction: (1) Sample background (prefer low-activity); (2) Sample antecedent $A$ and target $T$; (3) Insert $A$ followed immediately by $T$ with small adjacency gap.
@@
Construction: (1) Sample anchor $A$ and target $T$; (2) Sample $K\in\{1,2,3\}$ and direction; (3) Position anchor with room for $K$ targets; (4) Insert $K$ target needles in order.
@@
Construction: (1) Sample pure background from one regime; (2) Positive: insert 1 cross-regime needle (anomaly) + $N$ same-regime distractors; Negative: insert only same-regime distractors.

### v2

@@ 2 Background
Approaches such as Softprompting or Time-LLM (Jin et al., 2024), map the sequence of NN patch embeddings to the language space. However, this fidelity comes at a prohibitive computational cost ($\mathcal{O}(N^{2})$) either through the LLM’s extended context window or within the time series encoder, which must perform self-attention over the full patch sequence before projection (Wang et al., 2025). Consequently, these methods struggle to scale to larger context lengths due to computational cost and large memory requirements.
@@ B.7 Task Details
Construction: (1) Sample background window; (2) Select activity regime (sedentary/active); (3) Insert NN needles from regime activities not in background; (4) For positive samples, ask about an inserted activity; for negative, ask about a non-inserted activity from the same regime.
@@
Construction: (1) Sample two distinct activities AA and $B$; (2) Sample background excluding both; (3) Randomly assign temporal order (50/50 balance); (4) Insert both needles sequentially.
@@
Construction: (1) Sample background (prefer low-activity); (2) Sample antecedent AA and target $T$; (3) Insert AA followed immediately by $T$ with small adjacency gap.
@@
Construction: (1) Sample anchor AA and target $T$; (2) Sample $K\in\{1,2,3\}$ and direction; (3) Position anchor with room for $K$ targets; (4) Insert $K$ target needles in order.
@@
Construction: (1) Sample pure background from one regime; (2) Positive: insert 1 cross-regime needle (anomaly) + NN same-regime distractors; Negative: insert only same-regime distractors.